OpenStudy (anonymous):

how do i do this? W=T(s1-s2)-Q solve for s2

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

Distribute the T first to simplify.

OpenStudy (anonymous):

Then move every term that does not have s2 in it to the other side with the W, then you'll see what the last step is.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

is the finl nswer W-s1-Q=s2

OpenStudy (anonymous):

What happened to T?

OpenStudy (anonymous):

and Q will be positive when moved to the side with the W. Remember that when transposing terms, if you change sides, change signs.

OpenStudy (anonymous):

s2=Ts1+Q-W/T

OpenStudy (anonymous):

so positive T at the bottom

OpenStudy (anonymous):

That's really close, I think you just have opposite signs. I'll work through it again just to be sure.

OpenStudy (anonymous):

Actually, I think what you have is correct, but put parentheses around everything you are dividing by T like (Ts1-Q-W)/T, but I think the Q should also be negative.

OpenStudy (anonymous):

Here's another method: \[W=T(s_1-s_2)-Q\] \[W+Q=T(s_1-s_2)\] \[(W+Q)/T=s_1-s_2\] \[s_2=s_1-(W+Q)/T\]